OpenStudy (anonymous):

The value of x in \[(-\pi,\pi)\] which satisfies the equation \[\large 8^{(1+ \left|cosx \right|+\left|\cos^{2}x \right|+\left|\cos^{3}x \right|+.......\infty )}=4^{3}\] is ???

OpenStudy (anonymous):

ques. from series and sequences.

OpenStudy (experimentx):

8*8 = 4^3

OpenStudy (experimentx):

is that power??

OpenStudy (experimentx):

if that is power, then then the value of summation is 2 |cos x| <= 1, since this has to converge < 1 we know for G.S. 1/(1 - |cos x|) = 2 1 - |cos x| = 1/2 => x = pi/6 or 5pi/6

OpenStudy (experimentx):

sorry pi/3 and 2pi/3 ... 60 and 120